· Question 16

An investor is considering the following zero-coupon bond for her Income Preservation Portfolio: Face value: $1,000 Years left until maturity:10 years. Assuming that the YTM of this bond is 10.4%, its "price" (or DCF value) is closest to: ·

Question 17

You hold a zero-coupon bond with a $1,000 par value and 10 years left until maturity in your Income and Growth Portfolio. According to your financial advisor, the bond's current market price is $459. Based on this information, the yield to maturity on this bond is closest to: ·

Question 18

RVC Industries has 200,000,000 shares outstanding. The firm's expected earnings are $350,000,000. It will payout 30% of its earnings as follows: 75% in dividends and the remaining 25% will be spent on share repurchases. The firm's earnings growth rate is 7%, and its cost of equity capital is 12%. Based on this information, the firm's price per share is:
Answer #1

(16)

Price of Bond = PV of Face Value = Face Value/[(1+YTM)years] = 1000/[(1+0.104)10] = 1000/1.10410 = 1000/2.6896 = 371.8

(17)

YTM of zero coupon Bond = (Face Value/Present Value)1/years -1 = (1000/459)1/10-1 =2.1786490.1 -1= 1.08098-1 = 8.098%

(18)

Current year dividends = (Earnings*Payout)*Dividends out of payout = 350000000*0.3*0.75 = 78750000

Therefore, Current year dividend per share = 78750000/200000000 = 0.39375

Share Price = Next year dividend/(Cost of capital-growth rate)

Therefore, Share Price at the beginning of current year would be 0.39375/(0.12-0.07) = 7.875

Share repurchased = Total Amount spent on Repurchase/Price = [(Earnings*Payout)-Dividends]/Price

= [(350000000*0.3)-78750000]/7.875 = 3333333 shares

Therefore, Shares left = 200000000-3333333 = 196666667

Share Price = Next year dividend per share/(Cost of Capital-Growth Rate)

Next year dividend per share(assuming same payout but all 30% will be used for dividends and not for repurchase) = Total Dividends/Number of shares

= [(Earnings*Growth Rate)*Payout]/Number of shares

= [(350000000*1.07)*0.3]/196666667 = 0.57127

Therefore, Share Price = 0.57127/(0.12/0.07) = 0.57127/0.05 = 11.4254

(If this was helpful then please rate positively. Than You:)